A sailor strikes the side of his ship just below the surface of the sea. He hears the echo of the wave reflected from the ocean floor directly below 10.0 s later. How deep is the ocean at this point? (Note: Use the bulk modulus method to determine the speed of sound in this fluid, rather than using a tabluated value.)
_____ m
Answer #1
the echo time t = 10 s
speed of the water wave v = √K/ρ   ................ (1)
bulkmodulus of the water K = 2*109 N/m2
density of sea water ρ = 1.025*103 kg/m3
the velocity v = 2d/t   ................... (2)
compare eq (1) and (2) , we get
             2d/t = √K/ρ
dept of the ocean is
           d = (t / 2)√K/ρ
              = (10 s/2)√[(2*109 N/m2)/(1.025*103 kg/m3)]
              = 6.98*103 m
             = 6.98 km
              = 7 km (approximately)